OpenStudy (anonymous):

4z+1=2-3 what is the solution, I am sorry about before they gave me the wrong question.

OpenStudy (anonymous):

Solve for z. First, you want to isolate z so you minus 1 from both sides. After that, you divide both sides by 4. What do you get?

OpenStudy (anonymous):

not sure can we do it step by step

OpenStudy (anonymous):

Sure. First, \[4z+1-1=2-3-1\] What do you get?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Correct. Now the next step. \[{4z \over 4}={-2 \over 4}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

got the answer? :)

OpenStudy (anonymous):

no

OpenStudy (theviper):

4z+1=2-3 4z+1=-1 4z=-1-1 4z=-2 z=-2/4=-1/2

OpenStudy (anonymous):

so the answer is 1/2

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

that is -1/2 is the solution

OpenStudy (theviper):

yes

OpenStudy (anonymous):

thanks guys

OpenStudy (theviper):

http://www.purplemath.com/modules/solvelin.htm